OpenStudy (anonymous):

Four high school friends will all be attending the same university next year. There are 14 dormitories on campus. find the probability that at least 2 of the friends will be in the same dormitory

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

at least two in the same dorm is easiest to compute by computing the probability that they are all in different dorms, and subtracting that from 1

OpenStudy (anonymous):

I do not get that can you explain further please?

OpenStudy (anonymous):

I do not know how to find that probability. That is what I am struggling with

OpenStudy (misty1212):

maybe it is easier to do it directly he is in some dorm room. the probability that any one friend is in that same dorm room is \(\frac{1}{14}\)

OpenStudy (misty1212):

the probability that two are in that same dorm room is \(\frac{1}{14}\times \frac{1}{14}\)

OpenStudy (misty1212):

the probability that thre are in that same dorm room is \(\left(\frac{1}{14}\right)^3\)

OpenStudy (misty1212):

and the probability that all four are is \(\left(\frac{1}{14}\right)^4\)

OpenStudy (misty1212):

add those numbers up that should give you the answer

OpenStudy (misty1212):

\[\left(\frac{1}{14}\right)^2+\left(\frac{1}{14}\right)^3+\left(\frac{1}{14}\right)^4\]should do it

OpenStudy (anonymous):

Uh that's not the answer

OpenStudy (anonymous):

the answer is .37, but I do not know how to get to that answer.

OpenStudy (amistre64):

i wonder if a binomial probability works here

OpenStudy (amistre64):

p = 1/14, q = 13/14 just trying to work out what the structure would look like if it was, maybe? \[P(x)=\binom4xp^xq^{4-x}\]

OpenStudy (amistre64):

nah, that doesnt get you the answer that you say is correct so its most likely a bad thought as well

OpenStudy (amistre64):

how do we know the answer is .37?

OpenStudy (anonymous):

that's why I am asking.

OpenStudy (amistre64):

im saying, is the answer given to you in the back of a book, or did you find someone solution online ... or how did you determine that it has to be .37 to start with?

OpenStudy (anonymous):

the teacher posted answers for us to check with

OpenStudy (amistre64):

http://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.339416.html this one gets the same results that i did ... using binomial ill check others

OpenStudy (amistre64):

thats really the only one i can find using the google

OpenStudy (anonymous):

ah forget about that problem can you help me on another though?

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

An urn contains 10 red balls, 10 green balls, and 5 white balls. 5 balls are selected. In how many ways can 5 balls be drawn if at least 3 are green.

OpenStudy (amistre64):

does this thought make sense? (10 choose 3) green + (15 choose 2) not green or (10 choose 4) green + (15 choose 1) not green or (10 choose 5) green + (15 choose 0) not green

OpenStudy (amistre64):

the 15 choose 0 is not a good thought :) that would be 1, but we want none since 5 are already picked.

OpenStudy (amistre64):

and those are prolly best as multiplications, not additions ... then it would work better

OpenStudy (amistre64):

or brute it for proof gggrr gggrw gggww ggggr ggggw ggggg as long as order doesnt matter

OpenStudy (amistre64):

if order matters gggrr , 10 gggrw , 20 gggww , 10 ggggr, 5 ggggw , 5 ggggg, 1

OpenStudy (anonymous):

but the answer is 16002 according to the answers my teacher put up.